# Increasing Trust in Language Models through the Reuse of Verified Circuits

### v8

@@ 4.2 Investigating Five-Digit Addition
Ablation experiments targeting the nodes revealed that the model depends only on nodes located in nine token positions (Figs 6 and 7). Further ablation experiments show that for these nine token positions, the model uses 36 nodes in predictions. The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 5 and Table  1), providing insight into the specific computations performed at each node. For each answer digit $A_{n}$, using test questions corresponding to the ST8, ST9 and ST10 categories, we performed PCA on the nodes yielding interpretable results. Specifically nine “node and answer-digit” combinations (see Figs 8 and  9) showed strong clustering of the questions aligned to the ST8, ST9 and ST10 categories.
The algorithm predicts the first answer digit, A5, at position P11. A5, which is always 0 or 1, is the most challenging to predict as it may rely on a long carry one cascade (e.g. 55555+44445=100000). An accurate algorithm must compute this cascade using the nodes located in positions P8 to P11. As illustrated in Figure 7, these nodes attend to all digit pairs from D4 D’4 to D0 D’0. Additionally, the PCA data, as shown in Figures 8 and 9, suggest that these nodes produce tri-state outputs. After our first two algorithm hypothesises failed testing (see App. H and I), we discovered that the model utilizes a minimal set of “carry one” information, leading to the development of the TriCase quanta. The model performs $A_{n}.ST$ using bigrams (see App. F) to map two input tokens to one result token e.g. “6” + “7” = ST10. In positions P8 to P11, the model does $A_{n}.ST$ calculations on all digit pairs from D4 D’4 to D0 D’0.
An MLP layer can be thought of as a “key-value pair" memory (Meng et al., 2022; Geva et al., 2021) that can hold many bigrams and trigrams. We posit our MLP implements the TriAdd function using bigrams and trigrams to calculate $A_{n}.SV$ values from $A_{n}.ST$ values.
For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes and verified each node’s role using ablation intervention. Figure 3 diagrams the algorithm, with node details in App. J. The model adheres to all known constraints and achieves six nines accuracy. We concluded this model instance is well-understood, well-functioning, and hence known-good.
The 1-layer model uses 21 nodes to achieve two nines (99%) accuracy. This model uses 36 nodes (an increase of 71%) to achieve six nines accuracy.
@@
Table 2: All addition models studied implement our addition algorithm. The algorithm SA, SC and $ST$ sub-tasks all exist for each digit and in appropriate token positions. For a sample model, this map shows the subtask locations. Interestingly, here each SA sub-task is shared across two attention heads.
4.3 Training n-digit Addition models
To investigate whether this algorithm is used widely, we first trained seven 2-layer addition models with 5-, 6- and 10-digits, using different seeds, a different optimizer, and changing the answer format to include a sign token (e.g. 111111+222222=+0333333).
These seven models all have very low loss (e.g. 1.5e-8) and six nines accuracy. (Details in Tab.6.)
@@ 4.4 Investigating n-digit Addition models
We developed a declarative method to outline each sub-task. For instance, a sub-task may focus on question digits D2 and D’2, affect answer digit A3, influence S0 but not S1 complexity questions, and have specific PCA results and ablation tests. Using this declaration, we identified nodes performing these sub-tasks across all models.
An algorithm hypothesis, such as the one in section 4.2, is described by the required sub-tasks and their relationships. For example, our addition algorithm specifies that the model must execute $A_{n}.ST$ sub-tasks for each question digit before the “=” token. We created a framework for declaring n-digit algorithm hypotheses and testing them against the sub-tasks found in each model, mapping the results (see Table 2 for an example).
Our seven 2-layer addition models all implement our addition algorithm. Given their six nines accuracy and implementation of the same algorithm, we can confirm these models as known-good.
@@ 4.5 Training n-digit Mixed models
To explore reuse, we initialized untrained models with a known-good addition-only model, then trained them to perform both subtraction and addition. We call these “mixed" models.
Specifically, we trained seven larger (6- or 10-digit, 2- or 3-layer, 3- or 4-head) models after initializing them with the weights from a known-good 2-layer 3-head addition model. The first 2 layers, first 3 heads of the mixed model were initialized with the addition model weights. We trained the mixed model with 80% subtraction and 20% addition batches. We enriched the (infinite) training dataset with rare addition and subtraction edge cases.
Some models achieved six nines accuracy and the others five nines. Attempts to “freeze" the inserted attention heads and/or MLP layers by periodically copying the addition weights back into the mixed model every 100 training steps resulted in lower accuracy. (Refer App.M and Tab.6)
Figure 4: Each mixed model was initialized with the weights from a known-good addition model, then was trained on 80% subtraction and 20% addition batches. A sample log loss graph (final loss 8.0e-9) is shown.
@@ 4.6 Investigating n-digit Mixed models
Unlike addition, subtraction question answers can be either positive or negative. Similar looking positive-answer (e.g. 10009-10000=+000009) and negative-answer (e.g. 10009-20000=-009991) subtraction questions can give answers that differ at several digit positions. We posited that the model treats three distinct question classes “addition", “positive-answer subtraction" and “negative-answer subtraction" differently. Ablation of nodes showed that some nodes only help predict one question class, some help predict two classes and some all three classes. It also showed that the inserted addition nodes are heavily used and the majority become polysemantic, performing both addition and subtraction calculations. (Details in Tab.3.)
@@
Table 3: Mixed models re-use most inserted addition-model nodes. Many inserted nodes become polysemantic during training - performing addition, positive-answer subtraction and negative-answer subtraction sub-tasks simultaneously. For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
To investigate the subtraction algorithm, and paralleling addition algorithm subtasks, we defined subtraction-specific sub-tasks Base Diff and Borrow One (see Tab. 4) and complexity measures (see App. G and  N).
@@
Table 4: We define 2 “positive-answer subtraction" and 2 “negative-answer subtraction" sub-tasks that parallel the addition sub-tasks Base Add SA and Carry One SC
We found that the inserted nodes that performed SA in the addition model, perform SA, MD and ND in the mixed model. (Refer Table 5.) These three sub-tasks are similar in that each performs a mapping from 100 input cases (10 $D_{n}$x 10 $D^{\prime}_{n}$) to 10 output cases (0..9). The mixed model “upgraded" these nodes to be polysemantic during training. Similarly, some SC addition nodes became polysemantic and now process SC, MB and/or NB .
@@
An accurate subtraction model must answer “cascading borrow one" questions like 100000-000001=+0099999. We define the “essence of borrow one" subtask $MT$ used by both positive-answer (M) and negative answer (N) subtraction questions. $MT$ is like addition’s $ST$ subtask:
Paralleling addition’s “cascading carry one" $A_{n}.SV$ calculation, we define “cascading borrow one" calculation sub-tasks $A_{n}.MV$ for positive-answer subtraction and $A_{n}.NV$ for negative-answer subtraction. We posit the MLP implements TriAdd-like functions using bigrams and trigrams to calculate $A_{n}.MV$ and $A_{n}.NV$ values from $A_{n}.MT$ values.
@@ 4.7 Mixed model question class detection
We posit that there is a specific circuit to detect whether a question is in the S, M or N class. If the question operator is “+" then the class is S, but if the question operator is “-" then the model must calculate if $D>=D^{\prime}$ to distinguish between the M and N classes. For accuracy, it needs this answer by the “=" token to predict the first answer token (the answer sign) as “+" or “-".
$D>=D^{\prime}$ can be derived from the $A_{n}.MT$ data. Alternatively, this calculation could be a distinct circuit using a new sub-task we define as:
@@
For a 4 digit question, $D>=D^{\prime}$ can be calculated as A3.GT=1 or (A3.GT=0 and (A2.GT=1 or (A2.GT=0 and (A1.GT=1 or (A1.GT=0 and A0.GT=1))). Our test for GT is that ablation causes the answer to change sign. The $MT$ calculation is very similar.
We found $A_{n}.MT$ and $A_{n}.GT$ subtasks in all mixed models. Usually (but not always) both subtasks are calculated by the same node. (Table 2 shows an example.) As both the $MT$ and GT approaches are valid, models can learn valid algorithm sub-task implementations that differ per answer digit.
We further posit that a node that implements say A2.SA, A2.MD and A2.ND does not know whether it is dealing with a S, M or N question, so it outputs all three possible answers to the residual stream. Another node must calculate the S, M or N distinction likely by attending to the question operator (OPR) and the answer sign (SGN). We define attention sub-tasks OPR and SGN for these calculations. Table 2 shows that for each answer digit one attention head attends to both OPR and SGN. We believe these heads transfer sufficient data to the residual stream to allow the MLP layer(s) to calculate the question class, and so select the appropriate output from the polysemantic A2.SA/MD/ND node.
4.8 Mixed model summary
The 7 mixed experimental models achieve five or six nines accuracy and contain the same sub-tasks, implying a common algorithm. However, without fully understanding this algorithm, we can’t confirm these models as known-good.
@@ 5 Conclusion
We successfully trained and verified known-good 5-, 6- and 10-digit, 2-layer, 3-head addition models implementing the same algorithm with minor variations.
We demonstrated component reuse by integrating an existing addition model into a larger “mixed" model for both addition and subtraction, achieving six nines accuracy. This integration helped us understand the mixed model’s algorithm. The mixed model reuses most inserted addition nodes, upgrading many to become polysemantic - performing both addition and subtraction subtasks simultaneously.
Our work supports (Michaud et al., 2023)’s assertion that many natural prediction problems can be broken into finite "quanta" computations essential for loss minimization. It also aligns with the idea that understanding a network’s functionality involves identifying and comprehending its sub-quanta.
6 Limitations and Future Work
Our research suggests several future directions. A key goal is to complete reverse-engineering the mixed model’s algorithm, understanding polysemantic nodes and OPR+SGN node output representations in the residual stream, and how MLP layers use this to predict answer digits accurately.
We aim to develop a comprehensive, known-good model for n-digit addition, subtraction, and multiplication. Our approach also could be applied to create known-good models in logical reasoning and planning.
Using smaller models to enhance LLM trustworthiness and capabilities is an active research area, including model composition (Bansal et al., 2024), LM up-scaling which essentially emulates fine-tuning a large pre-trained model using a small model (Mitchell et al., 2023), and inserting accurate models into untrained ones. Research on “spare” neurons in LLMs ((Voita et al., 2023; Hu et al., 2021)) suggests potential for small-scale modifications to fix erroneous circuits. Further exploration of using known-good models to improve LLMs is warranted.
Developing methods to incorporate compact known-good models into LLMs could democratize AI Safety research, allowing small teams to focus on specific areas and create quality components to improve LLMs.
Our automated framework for discovering algorithm sub-tasks in models accelerated our research and can test whether known-good components retain functionality when integrated into LLMs. While some aspects are specific to our math models, much is generic and reusable. Extending the framework to integrate output from complementary tools (e.g. ACDC’s node connectivity strength data (Conmy et al., 2023)) would provide more automated insights.
We progressed on a declarative language to describe algorithms in terms of necessary sub-tasks and a framework to test these descriptions against specific models. Further work is needed to fully realize its potential.
7 Impact Statement
@@ Appendix A Appendix: Terminology
  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+" or "-" answer sign
@@
  • SA : Basic Add. An addition sub-task. An.SA is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
  • SC : Carry One. An addition sub-task. An.SC is defined as Dn + D’n >=>=> = 10. e.g. 5 + 7 gives True
  • SS : Make Sum 9. An addition sub-task. An.SS is defined as Dn + D’n == 9. e.g. 5 + 7 gives False
  • ST : TriCase. An addition sub-task. Refer paper 2 for details
  • ST8, ST9, ST10 : Outputs of the ST TriCase sub-task.
@@
  • : Basic Difference. A subtraction sub-task. An.MD is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • : Borrow One. A positive-answer subtraction sub-task. An.MB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • MZ : Make Zero. A positive-answer subtraction sub-task. An.MZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • MT : TriCase. A positive-answer subtraction sub-task.
  • MT1, MT0, MT-1 : Outputs of the MT TriCase sub-task.
@@
  • ND : Basic Difference. A negative-answer subtraction sub-task. An.ND is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • NB : Borrow One. A negative-answer subtraction sub-task. An.NB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • NZ : Make Zero. A negative-answer subtraction sub-task. An.NZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • NT : TriCase. A negative-answer subtraction sub-task.
  • GT : Greater Than. A (positive-answer or negative-answer) subtraction sub-task. An.GT is defined as Dn $>$ D’n. e.g. 3 $>$ 5 gives False
  • OPR : Operator. A sub-task that attends to the + or - token in the question (which determines whether the question is addition or subtraction).
  • SGN : Sign. A sub-task that attends to the first answer token, which is + or -

### v9

@@ 4.2 Investigating Five-Digit Addition
Through ablation experiments, we found that the model relies on just 36 nodes distributed across nine token positions (Figs 7 and 8). We analyzed how these nodes affect computation complexity and answer accuracy through ablation studies (see Figs 6 and Table 1).
Core Algorithm Discovery. The model’s most challenging task is predicting A5, the first answer digit, at position P11. This digit, which can only be 0 or 1, depends on potential carry-bit cascades through multiple positions (e.g., in 55555+44445=100000). After two unsuccessful algorithm hypotheses (detailed in App.H andI), we discovered the model’s fundamental mechanism: it uses a minimal representation of carry information through what we call the TriCase quanta. The model computes $A_{n}.ST$ using bigram operations that map digit pairs to result tokens (e.g., ”6” + ”7” = ST10) in positions P8 to P11.
For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes and verified each node’s role using ablation intervention. Figure 3 diagrams the algorithm, with node details in App. J. The model adheres to all known constraints and achieves >99.999% accuracy using 36 nodes, as compared to the 1-layer model uses 21 nodes to achieve 99% accuracy.
Table 2: All addition models studied implement our addition algorithm. The algorithm SA, SC and $ST$ subtasks all exist for each digit and in appropriate token positions. For a sample model, this map shows the subtask locations. Interestingly, here each SA subtask is shared across two attention heads.
@@ 4.3 Investigating n-digit Addition models
To test whether this algorithm generalises beyond 5-digit addition, we conducted a comprehensive study of larger addition models. We trained 15 2-layer transformer models to handle addition problems ranging from 5- to 15-digits. These models used varying random seeds and optimizers, and included a sign token in their output format (e.g. 111111+222222=+0333333).
Figure 4: The 5- to 13-digit 2-layer 3-head addition models have very low loss and >99.999% accuracy. With more digits the model loss and accuracy is worse. Details in Tab.6.
Analysis Framework. To systematically analyse these models, we first created a declarative method for specifying each subtask within the addition process. Each subtask definition includes (1) the input digits it processes, (2) the answer digit it affects, (3) the complexity level it influences, and (4) expected PCA patterns and ablation test results. Using this declaration, we identified the nodes performing these subtasks across all models.
Since our framework lays out the addition algorithm in terms of its required subtasks., we can also verify whether a given model implements this algorithm by checking for the presence and proper sequencing of all required subtasks. For example, our addition algorithm specifies that the model must execute $A_{n}.ST$ subtasks for each question digit before the first answer token. Table 2 shows how an addition model fully implements all required subtasks in its nodes.
Using this framework, we found that all fifteen 2-layer addition models, regardless of their size, implemented the same core algorithm we identified in our 5-digit model. This consistent pattern suggests we have identified a robust and generalizable algorithm that naturally arises when transformer models are trained on addition in our setup where the digits are tokenized individually.
@@ 4.4 Training n-digit Mixed models
To explore how transformers can learn multiple arithmetic operations simultaneously, we conducted experiments with models capable of both addition and subtraction. We were particularly interested in whether knowledge from addition-specialized models could transfer to this expanded task.
We trained 14 mixed models with various architectures, including 6- and 10-digit models with 2-3 layers and 3-4 attention heads. Each model was initialized using weights from a highly accurate addition-only model, specifically copying the weights into the first two layers and first three heads of the mixed model. During training, we used a curriculum of 80% subtraction and 20% addition examples, with deliberate enrichment of rare edge cases for both operations.
This initialization approach proved effective, with most models achieving high accuracy on both operations. We explored alternative training strategies, such as periodically ”freezing” the inserted attention heads or MLPs by restoring their original addition-model weights every 100 training steps. However, these more constrained approaches resulted in lower accuracy, suggesting that some flexibility in modifying the initialized weights was necessary for optimal performance.
The initialized mixed experimental models achieve >99.999% accuracy and contain the same subtasks, implying a common algorithm.
Figure 5: The 5- to 8-digit initialized mixed models have very low loss and >99.999% accuracy. With more digits the model loss and accuracy is worse. Details in Tab.7.
@@ 4.5 Investigating n-digit Mixed models
The mixed models revealed fascinating patterns in how they handled different types of arithmetic problems. Our hypothesis was that the model handled three distinct question classes: addition, positive-answer subtraction (e.g., 10009-10000=+000009), and negative-answer subtraction (e.g., 10009-20000=-009991).
Through ablation studies, we found that many nodes specialized in handling specific question classes, while others became ”polysemantic,” capable of processing multiple classes simultaneously (More details in Tab.3). Most notably, the majority of nodes initialized from the addition model became polysemantic during training. Rather than developing entirely new circuits for subtraction, the models adapted their addition circuits to handle both operations.
Table 3: Mixed models re-use most inserted addition-model nodes. Many inserted nodes become polysemantic during training - performing addition, positive-answer subtraction and negative-answer subtraction subtasks simultaneously. For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
@@
Defining Subtraction Subtasks. Subtraction requires two distinct computational paths based on the relative magnitude of the numbers. When D $\geq$ D’, as in 325 - 123 = +202, we perform direct subtraction with borrowing. When D $<$ D’, as in 325 - 456 = -131, we must instead compute the magnitude of the difference and mark it negative.
Hence, this computational difference necessitates separate frameworks for positive-answer (M) and negative-answer (N) subtraction, each with parallel operations to addition’s Base Add ($SA$) and Carry One ($SC$), which we define in Tab. 4. We also define complexity measures for subtraction (see App. G and  N).
Table 4: We define 2 “positive-answer subtraction” and 2 “negative-answer subtraction” subtasks that parallel the addition subtasks, Base Add SA and Carry One SC .
@@
Just as addition must handle ”cascading carry” cases (e.g., 99999+1=100000), subtraction must handle ”cascading borrow” cases. Consider 100000-1=099999, where a borrow cascades through multiple zeros. To handle such cases efficiently, we define a three-state classification $MT$ that parallels addition’s $ST$ :
Finally, paralleling the ”cascading carry one” $An.SV$ operation in addition, we define cascading borrow calculations $A_{n}.MV$ and $A_{n}.NV$ for positive and negative cases.
Evolution of Polysemantic Nodes. During training, the mixed model adapted the nodes it inherited from the addition model. Nodes that originally only performed Simple Addition ($SA$) developed the ability to simultaneously handle three distinct operations: addition ($SA$), positive-answer subtraction ($MD$), and negative-answer subtraction ($ND$). (Refer to Table 5.) We believe that this adaptation was likely because these three operations share a common structure: each maps a pair of digits (10 possibilities for $D_{n}$× 10 possibilities for $D^{\prime}_{n}$= 100 input cases) to a single-digit result (10 possible outputs). The model similarly adapted nodes that originally handled carry bits ($SC$) to also process borrow operations for both positive ($MB$) and negative ($NB$) subtraction cases.
@@ 4.6 Mixed model question class detection
Mixed models must identify whether each input belongs to one of three classes: addition (S), positive-answer subtraction (M), or negative-answer subtraction (N). For addition questions (marked by ”+”), classification is immediate. For subtraction questions (marked by ”-”), the model must determine whether D $\geq$ D’ to distinguish between M and N classes. This classification must occur by the ”=” token since it determines whether the first answer token will be ”+” or ”-”.
The model can determine D $\geq$ D’ in two ways. First, it can use the $A_{n}.MT$ values already computed for subtraction. Alternatively, it can use a dedicated comparison circuit based on a new subtask:
@@
For a 4-digit number, this comparison cascades through the digits: D $\geq$ D’ is true if either $A3.GT=1$, or if $A3.GT=0$ and $A2.GT=1$, and so on through $A0.GT$. We verify this circuit through ablation tests that show its disruption causes answer signs to flip.
We found all mixed models incorporate the $A_{n}.MT$ and $A_{n}.GT$ subtasks to determine whether the answer is postiive or negative. Typically (but not always), the same node computes both subtasks (see Table 2). Since both the $MT$ and GT approaches can validly distinguish between positive and negative outcomes, the model sometimes chooses one approach for some digits and the other approach for others.
A single node can implement multiple arithmetic operations (e.g., $A2.SA$, $A2.MD$, $A2.ND$) but has no intrinsic way of knowing whether the current question is addition, positive subtraction, or negative subtraction. Hence, we hypothesise that it outputs all three possible results into the residual stream. A different node must then select the correct result by identifying the question’s operation and required sign.
In practice, we observe that one or more attention heads attend to the question operator token (OPR) and the initial answer sign token (SGN), encoding the necessary information for the downstream MLP layers to pick which of the three outputs is correct. As Table 2 shows, each answer digit typically has one attention head that focuses on both the OPR and SGN tokens, thereby enabling the MLP to filter out the irrelevant outputs and retain the correct one.
@@ 5 Conclusion
Our analysis of 44 transformer models demonstrates their ability to learn highly accurate mathematical algorithms, achieving over 99.999% accuracy on addition tasks up to 13 digits. By investigating how models implement computations using attention heads and MLPs, we uncovered novel computational approaches distinct from traditional human calculation methods. Parameter transfer experiments revealed that features originally specialized for addition could be repurposed for subtraction, with many nodes becoming polysemantic.
Our research provides insights into how transformers learn precise algorithmic computations across related tasks, developing a reusable library of mechanistic interpretability tools that advances our understanding of neural network computational learning.
Impact Statement
@@ Appendix A Appendix: Terminology
  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+” or ”-” answer sign
@@
  • SA : Basic Add. An addition subtask. An.SA is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
  • SC : Carry One. An addition subtask. An.SC is defined as Dn + D’n >=>=> = 10. e.g. 5 + 7 gives True
  • SS : Make Sum 9. An addition subtask. An.SS is defined as Dn + D’n == 9. e.g. 5 + 7 gives False
  • ST : TriCase. An addition subtask. Refer paper 2 for details
  • ST8, ST9, ST10 : Outputs of the ST TriCase subtask.
@@
  • : Basic Difference. A subtraction subtask. An.MD is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • : Borrow One. A positive-answer subtraction subtask. An.MB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • MZ : Make Zero. A positive-answer subtraction subtask. An.MZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • MT : TriCase. A positive-answer subtraction subtask.
  • MT1, MT0, MT-1 : Outputs of the MT TriCase subtask.
@@
  • ND : Basic Difference. A negative-answer subtraction subtask. An.ND is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • NB : Borrow One. A negative-answer subtraction subtask. An.NB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • NZ : Make Zero. A negative-answer subtraction subtask. An.NZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • NT : TriCase. A negative-answer subtraction subtask.
  • GT : Greater Than. A (positive-answer or negative-answer) subtraction subtask. An.GT is defined as Dn $>$ D’n. e.g. 3 $>$ 5 gives False
  • OPR : Operator. A subtask that attends to the + or - token in the question (which determines whether the question is addition or subtraction).
  • SGN : Sign. A subtask that attends to the first answer token, which is + or -